OpenStudy (anonymous):

solve for the given domain. write the solution as a set of ordered pairs.2x-2y=3, domain{-1, 0,1,2}

OpenStudy (ivettef365):

just find out what is y when x = -1,0,1,2

OpenStudy (anonymous):

i still dont get that. so x= all the numbers?

OpenStudy (ivettef365):

ok, you have for x=-1 so when x = -1 lets find out what is y 2x -2y = 3 2(-1) -2y = 3 -2 -2y = 3 -2y = 3 + 2 -2y = 5 y = -5/2

OpenStudy (ivettef365):

now lets find out when x = 0 what is y 2x -2y = 3 2(0) -2y = 3 -2y = 3 y = -3/2

OpenStudy (ivettef365):

you have 2 more to go

OpenStudy (anonymous):

okay. so then when i get those, how do i find out the orderd pairs?

OpenStudy (ivettef365):

ordered pair (x,y)

OpenStudy (ivettef365):

so you have (-1, -5/2)

OpenStudy (anonymous):

ohkay thanks soo much i get it now. i remeber learning this